OpenStudy (surryyy):

Which values for x and y make the statement (x + 4)(y - 1) = 0 true? x = -4, y = 1 x = -4, y = -1 x = 4, y = 1 x = 4, y = -1

imqwerty (imqwerty):

the product of 2 numbers is 0 if either of the two is 0 or if both are zero so we have - (x + 4)(y - 1) the product can be zero if x+4 becomes 0 so x+4= 0 so x=-4 the product can become 0 when y-1=0 so y=1 so 1st option is correct :)

OpenStudy (surryyy):

oh ok tysm!

imqwerty (imqwerty):

welcome :D

OpenStudy (surryyy):

Which values for x and y make the statement (x - 3)(y + 8) = 0 true? x = -3, y = 8 x = 3, y = 8 x = 3, y = -8 x = -3, y = -8 for this one would my answer be c?

imqwerty (imqwerty):

yes :)

OpenStudy (surryyy):

ok thnx!

imqwerty (imqwerty):

no problem :D